A block of mass 1kg moving with a speed 4m/s collides with another block of mass 2kg at rest,the light block comes to rest after collision,the loss in k.e. of the system is

Grade:11

1 Answers

vikas askiitian expert
13 years ago

m1u1=m1v1 + m2v2   (from conservation of momentam)

 1*4 =  2*v2 (v1=0 ..given)

 v2=2m/s

 (kE)i =m1(u1)^2 /2

         =8joule

 (kE)f=m2(v2)^2 /2

       =4j

 so kinetic energy transfered is 50%